a 10% and a 20% salt solution are combined to make 12 gallons of 15% salt solution. How much of each strength solution is required?

Answer by mananth(16946)

percent ----------------quantity
salt solution I 10---------------x lb
salt solution II20---------------12-x lb
Total 15.00% ---------------- 12

10x+20(12-x)=15 *12
10x+240-20x=180
10x-20x=180-240
-10x=-60
/-10
x=6lb ---10.00%
6lb---- 20.00%